# Vulnerabilities in AI-generated Image Detection: The Challenge of Adversarial Attacks

### v1

@@ 4.3.1. Attack SOTA AIGI Detectors
Table 6. The attack success rate (ASR) is reported on real images and fake images respectively. The surrogate model is chosen as Swin-Vit.
@@
4.3.2. Attack Cross-Generator Image Detection
@@
One important real-world detection problem is cross-generator image detection, i.e. identifying fake images generated by unseen generative models. We hence evaluate the robustness of cross-generator image detection to investigate whether adversarial examples are a real threat to AIGI detection. We train the detectors on images generated by SD v1.4 (Rombach et al., 2022) and assess their robustness against adversarial examples, in which the adversarial perturbations are added to the images generated by Midjourney (Midjourney, 2022), SD V1.4 (Rombach et al., 2022), SD V1.5 (Rombach et al., 2022), ADM (Dhariwal and Nichol, 2021), Wukong (Wukong, 2022) and BigGAN (Brock et al., 2018). Because Swin-Vit achieves the SOTA results on different subsets (Zhu et al., 2024), we use it as the surrogate model.
@@
The benign accuracy and attack performance on unseen source data are reported in Tab. 4 and Tab. 5 respectively. First, the attack under the white-box setting achieves almost 100% success rate. Second, the transfer success rate is positively correlated to the accuracy on unseen source data. The Midjourney, SD v1.4&v1.5 and Wukong subsets have relatively high accuracy, their corresponding transfer attack success is also relatively high. In contrast, the binary classification accuracy drops to 50% on ADM and BigGAN subsets, the corresponding adversarial transferability is also limited on them. By looking closely at the accuracy/ASR on real images and fake images(Tab. 4,Tab. 6), we find detectors fail to distinguish fake images on ADM and BigGAN, which means without a good gradient can be followed to misclassify the real image as fake label under the attack. Therefore, we suggest robustness evaluation of cross-generator image detection should be conducted on the test subset with high accuracy as evaluating on the low-accuracy subset is futile.
4.3.3. Attack Defense Models
Furthermore, we evaluate FPBA on attacking defense models. We first employ adversarial training (Madry et al., 2017) for evaluation, which is the most popular defense baseline. However, according to our preliminary experiments(reported in the supplementary material), the training process of adversarial training fails to converge, which shows adversarial training is not a reliable way for evaluating attacks on AIGI detection. Considering that AIGI detection typically (Wang et al., 2020) uses JPEG compression and Gaussian blurring as data preprocessing during training to improve their robustness, we follow the suggestions from (Wang et al., 2020; Zhu et al., 2024) to utilize Gaussian blurring and JPEG compression as a defense. As shown in Tab. 7, existing attack methods all suffer from performance degradation as the probability of using Gaussian blurring and JPEG compression increases during training, while FPBA still contains a high attack success rate, which validates its effectiveness.
@@ 4.3.3. Attack Defense Models
Table 7. Defense models(Blur+JPEG) against adversarial attack on ProGAN Dataset. Blur+JPEG(0.1) and Blur+JPEG(0.5) represented that images are blurred and JPEG-ed with 10% and 50% probability respectively. We report the attack success rate(%).
@@
SSA 100.0 80.9(-19.1)
FPBA 100.0 99.2(-0.5)
4.4. Additional Performance Analysis
4.4.1. The Phenomenon of Gradient Masking in AIGI Detectors
In Tab. 1, we find that the attack results of MIFGSM are similar to IFGSM, and significantly lower than PGD, in contrast to the common belief that adversarial examples generated by momentum iterative methods have higher success rate (Dong et al., 2018). It indicates the possibility of gradient masking (Athalye et al., 2018). To investigate whether AIGI detectors exist in gradient masking, we analyze the aggregated gradient convergence properties. We take the adversarial examples crafted by ResNet-50 on ProGan as an example. For each plot, we randomly sample 500 adversarial examples generated by a specific attack to compute their expected loss gradients. Each dot shown in Fig. 6 represents a component of the expected loss gradient from each image, in which there are a total of 75k loss gradient components.
Figure 6. The gradient components of ResNet on adversarial examples generated by different attack methods. (a) gradient components of IFGSM; (b) gradient components of PGD; (c) gradient components of MIFGSM; (d) gradient components of MIFGSM with random initialization.
4.4.2. Qualitative Analysis
Table 8. The visual quality of different attack samples in terms of the average MSE, PSNR and SSIM scores.
@@ 4.4.2. Qualitative Analysis
Figure 7. Visual comparison with PGD. (a): the original image generated by diffusion models. (b): adversarial examples crafted by PGD. (c): adversarial examples crafted by FPBA. The image quality of the adversarial example crafted by our method is much closer to the original image.
In this study, we evaluated the image quality of the generated adversarial samples using PSNR and SSIM scores as metrics. We present the quantitative results of MSE, PSNR(db) and SSIM in Tab. 8. The image quality of our method is higher than PGD. As illustrated in Fig. 7, FPBA can generate more natural-looking adversarial examples in comparison with PGD.
4.4.3. Ablation Study
Table 9. Ablation Study on ProGAN dataset. The adversarial samples are crafted from ResNet50.
Attack DenNet MobNet Spec ViT
@@ 4.4.3. Ablation Study
We conduct an ablation study in Tab. 9 to investigate the impact of computing gradient in different domains. In comparison with computing attack gradient solely in the spatial domain or frequency domain, our spatial-frequency attack achieves higher transferability. More ablation studies can be found in the supplementary material.
5. Conclusion
@@
These findings indicate that building robust AIGI detectors is still an open problem. We hope that our work can be used as a protocol to evaluate the robustness of AIGI detectors and draw more attention to the robustness of AIGI detectors.

### v3

@@ IV-B3 Evaluation on Frequency-based Detectors
TABLE VII: The attack success rate (ASR) is reported on real images and fake images respectively. The surrogate model is chosen as Swin-ViT.
@@
IV-C Evaluation on Diverse Detection Strategies
IV-C1 Attack Various SOTA Detection Methods
Except for detecting in spatial and frequency domian using different backbone models, recent AIGI detectors utilize different latent feature representations to identify synthetic artifacts. Specifically, GramNet [53] extracts the global texture representation; UnivFD [24] captures semantic feature using pre-trained CLIP [58]; LNP [22] extracts the noise pattern of images; LGrad [23] extract gradient information through a pre-trained model and DNF [28] leverages diffusion noise features, extracting them through an inverse diffusion process [2]. These significant differences among various detection methods present challenges for the transferability of adversarial examples. To illustrate this, we examine the adversarial transferability of various detection methods and present the results in Table IV. Overall, the transfer success rate is relatively low compared to the results in Table I, suggesting that the distinct characteristics of detection methods limit adversarial transferability on each other. Nevertheless, FPBA still give the best adversarial transferability.
IV-C2 Attack Cross-Generator Image Detection
@@
One important real-world detection problem is cross-generator image detection, i.e. identifying fake images generated by unseen generative models. We hence evaluate the robustness of cross-generator image detection to investigate whether adversarial examples are a real threat to AIGI detection. We train the detectors on images generated by SD v1.4 [45] and assess their robustness against adversarial examples, in which the adversarial perturbations are added to the images generated by Midjourney [59], SD V1.4 [45], SD V1.5 [45], ADM [60], Wukong [61] and BigGAN [62]. Because Swin-ViT achieves the SOTA results on different subsets [12], we use it as the surrogate model.
@@
The benign accuracy and attack performance on unseen source data are reported in Table V and Table VI respectively. First, the attack under the white-box setting achieves almost 100% success rate. Second, the transfer success rate is positively correlated to the accuracy on unseen source data. The Midjourney, SD v1.4&v1.5 and Wukong subsets have relatively high accuracy, their corresponding transfer attack success is also relatively high. In contrast, the binary classification accuracy drops to 50% on ADM and BigGAN subsets, the corresponding adversarial transferability is also limited on them. By looking closely at the accuracy/ASR on real images and fake images(Table V,Table VII), we find detectors fail to distinguish fake images on ADM and BigGAN, which means without a good gradient can be followed to misclassify the real image as fake label under the attack. Therefore, we suggest robustness evaluation of cross-generator image detection should be conducted on the test subset with high accuracy as evaluating on the low-accuracy subset is futile.
IV-D Attack Performance against Compression in Real Scenarios
Figure 6: Attack performace on ResNet, Efficient, MobileNet and Swin against JPEG compression on LSUN(ProGAN) dataset. Adversarial samples are crafted from ResNet.
In real-world scenarios, images are often compressed during propagation, which may also distort adversarial noise patterns. Therefore, ensuring the effectiveness of adversarial attacks on compressed images is crucial for simulating real-world scenarios. To this end, we perform JPEG compression with quality factors of 30, 40, 50, 60, 70, 80, 90 and 100 to the tested images. We evaluate the degradation in attack performance using four models: ResNet-50, EfficientNet, MobileNet and Swin Transformer. All adversarial samples are crafted from ResNet-50. The results are reported in Fig. 6. For white-box attacks on ResNet, JPEG compression has a negligible effect on the attack success rate across different attack methods, maintaining top-1 attack performance. For transfer-based attacks, the attack strength on victim models decreases as the quality factor decreases. However, FPBA consistently maintains the highest attack success rate, demonstrating the effectiveness of our attack method under the real-world scenarios.
@@ IV-D Attack Performance against Compression in Real Scenarios
TABLE VIII: Defense models(Blur+JPEG) against adversarial attack on ProGAN Dataset. Blur+JPEG(0.1) and Blur+JPEG(0.5) represented that images are blurred and JPEG-ed with 10% and 50% probability respectively. We report the attack success rate(%).
@@
S2I 100.0 80.9(-19.1)
FPBA 100.0 99.2(-0.8)
IV-D1 Attack Defense Models
Furthermore, we evaluate FPBA on attacking defense models. We first employ adversarial training [38] for evaluation, which is the most popular defense baseline. However, according to our preliminary experiments(reported in the supplementary material), the training process of adversarial training fails to converge, which shows adversarial training is not a reliable way for evaluating attacks on AIGI detection. Considering that AIGI detection typically [3] uses JPEG compression and Gaussian blurring as data preprocessing during training to improve their robustness, we follow the suggestions from [3, 12] to utilize Gaussian blurring and JPEG compression as a defense. As shown in Table VIII, existing attack methods all suffer from performance degradation as the probability of using Gaussian blurring and JPEG compression increases during training, while FPBA still contains a high attack success rate, which validates its effectiveness.
TABLE IX: The attack success rate(%) on CNN-based, ViT-based and Frequency-based models on the Synthetic FFHQ datasets. “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model. We mark the white-box attack results in gray, and black-box attack results are not marked with colors.
Surrogate Model Attack Methods CNNSpot DenseNet EfficientNet MobileNet Spec DCTA ViT Swin Average
IFGSM 67.5 67.5 55.3 67.5 28.9 32.5 16.2 17.5 40.8
MIFGSM 67.5 67.5 54.9 67.6 28.8 32.6 20.7 17.9 41.4
PGD 100 100 88.6 99.9 49.9 50 40.7 50.0 68.4
CNNSpot S2I 100 100 77.7 90.1 46.7 52.9 37.6 55.8 65.8
SSAH 99.0 54.5 44.1 41.0 2.6 7.2 1.0 4.2 22.1
FPBA(Ours) 100 100 86.8 100 49.4 49.6 43.1 50.9 68.5
IFGSM 10.7 12.7 23.8 60.4 24.5 28.4 7.3 10.2 16.8
MIFGSM 10.3 12.6 23.2 60.4 25.4 28.7 9.1 10.2 17.1
PGD 49.8 50.2 57.5 100 49.7 49.8 38.8 49.8 49.4
MobileNet S2I 82.8 86.7 78.5 100 51.4 51.8 28.9 49.7 61.4
SSAH 1.9 36.4 43.7 98.5 2.0 6.2 0.8 2.4 13.3
FFHQ (Style GAN) FPBA(Ours) 51.2 55.2 66.6 100 50.3 49.9 39.4 49.8 51.8
IV-E Visual Quality Analysis
TABLE X: The visual quality of different attack samples in terms of the average MSE, PSNR and SSIM scores.
@@ IV-E Visual Quality Analysis
Figure 7: Visual comparison with PGD. (a): the original image generated by diffusion models. (b): adversarial examples crafted by PGD. (c): the original image generated by FakePolisher. (d): adversarial examples crafted by FPBA. The image quality of the adversarial example crafted by our method is much closer to the original image.
To demonstrate the superior image quality achieved by our method, we conduct both qualitative and quantitative assessments. We first visualize adversarial examples generated by PGD, FakePolisher and our proposed FPBA in Fig. 7. Adversarial examples from PGD and FakePolisher show noticeable noise patterns upon zooming in. In addition, the fine-gradined details of object generated by FakePolisher are blurred, which is very obvious in the gota’s texture and outline in its adversarial images. In contrast, FPBA generate more natural-looking adversarial examples. Further, we also report the quantitative results using common metrics for image quality assessment including MSE, PSNR(db) and SSIM. As reported in Table X, FPBA outperforms other baselines across all quality assessment by a large margin. This suggests that FPBA adding adversarial noise in the frequency-domian, rather than directly in the spatial domian, is more imperceptible to observers.
IV-F Additional Performance Analysis
IV-F1 The Phenomenon of Gradient Masking in AIGI Detectors
In Table I, we find that the attack results of MIFGSM are similar to IFGSM, and significantly lower than PGD, in contrast to the common belief that adversarial examples generated by momentum iterative methods have higher success rate [54]. It indicates the possibility of gradient masking [63]. To investigate whether AIGI detectors exist in gradient masking, we analyze the aggregated gradient convergence properties. We take the adversarial examples crafted by CNNSpot on ProGan as an example. For each plot, we randomly sample 500 adversarial examples generated by a specific attack to compute their expected loss gradients. Each dot shown in Fig. 8 represents a component of the expected loss gradient from each image, in which there are a total of 75k loss gradient components.
Figure 8: The gradient components of CNNSpot on adversarial examples generated by different attack methods. (a) gradient components of IFGSM; (b) gradient components of PGD; (c) gradient components of MIFGSM; (d) gradient components of MIFGSM with random initialization.
IV-F2 Evaluation on Deepfake Datasets
As we mention before, there is a big difference between AIGI detection and Deepfake detection. But we empirically found that FPBA is also effective for face forgery detection. To demonstrate this, we conduct the experiments on synthetic FFHQ face datasets [46], which is consist of 50k real face images from FFHQ and 50k fake generated face images created by StyleGAN2 [47]. We report the results in Table IX. Although FPBA is not specific designed for Deepfake detection, it still achieve the best white-box attack performance. For transfer-based attack, FPBA also has the top-2 performance, demonstrating that FPBA is a universal threat across both AIGI and Deepfake detection.
TABLE XI: Ablation Study on ProGAN dataset. The adversarial samples are crafted from CNNSpot.
Attack DenseNet MobileNet Spec ViT
@@ IV-F2 Evaluation on Deepfake Datasets
TABLE XII: The attack success rates (%) of FPBA on normally trained detectors w.r.t the number $N$ of spectrum transformations. “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model.
$N$ ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Ave.
0 64.5 64.0 37.3 60.6 17.5 48.4 40.5 59.7 48.8
5 98.9 98.0 76.4 95.9 19.8 48.0 51.5 94.8 69.2
10 99.6 99.1 78.6 97.5 19.9 48.0 52.4 96.5 70.3
15 99.4 99.0 78.2 96.6 20.2 48.3 53.0 96.2 70.2
20 99.2 98.7 77.5 96.7 20.7 48.6 51.1 95.6 69.8
IV-G Ablation Study
Spatial and Frequency Attack
We conduct an ablation study in Table XI to investigate the impact of computing gradient in different domains. In comparison with computing attack gradient solely in the spatial domain or frequency domain, our spatial-frequency attack achieves higher transferability. More ablation studies can be found in the supplementary material.
Number ($N$) of Spectrum Transformations
In this study, we investigate the impacts on the number ($N$) of spectrum transformations, which can simulate diverse substitute models. The adversarial examples were crafted from ResNet-50 in LSUN(ProGAN) dataset. With the exception of $N$, other hyper-parameters keep the same with the default settings in the paper. As shown in Table XII, with $N$ increasing 5 from 0, FPBA shows a substantial augmentation in attack success rate. Further, when $N>5$, there is a diminishing gain in attack success rate but with increased computation. Considering the trade-off between attack performance and computation cost, we set $N=5$ by default.
V Conclusion
@@
These findings indicate that building an adversarial robust AIGI detector is still an open problem. We hope that our work can be used as a protocol to evaluate the robustness of AIGI detectors and draw more attention to this domain.
Acknowledgments
This work was supported in part by the NSF China (No. 62302139) and FRFCU-HFUT (JZ2023HGTA0202, JZ2023HGQA0101).